# The emergence of galactic thin and thick discs across cosmic history

### v1

@@ 3.3 Emerging thin discs at later epochs of universe
Figure 9 shows the distribution of two-disc and single-disc galaxies in stellar mass and lookback time, along with fractions of two-disc galaxies as a function of each variable. Two-disc galaxies are prevalently found at high stellar masses at earlier times, extending back to $\sim$10 Gyr ago ($z\sim 2$). At later times (i.e., shorter lookback times), they are increasingly found at lower stellar masses, resulting in a increasing fraction of two-disc galaxies with cosmic time and stellar mass. The distribution includes observational biases: identifying a two-disk structure becomes more difficult for more distant galaxies or lower mass galaxies. To dissect observational effects from the onset of thin disc formation, Fig. 9 shows the range of lookback times at which two disks can be identified when a galaxy (of similiar properites) with two discs is artificially redshifted (gray line for each data point).
To find the maximum lookback time, we calculate redshifted surface brightness, disc radius ($h_{R}$), and disc height ($z_{0}$) by redshift increments ($dz=0.02$) using galaxies’ best-fit model parameters for galaixes categorised as having two discs (Sec. 3). We consider the surface brightness evolution including not only the surface brightness dimming ($\propto(1+z)^{-4}$) but also k-correction and intrinsic evolution (see Fig. 25 in Appendix E). We then generate simulated images with noise added and the identical PSF, and refit the images using both a two-disc model and one-disc model. The noise incorporates variance from sky noise, read noise, and Poisson variance due to photon counting from the object itself. Bulge components are subject to redshifting but fixed during the refit. We consider the two-discs to be detectable if the BIC difference between the two-disc fit and one-disc fit is conservatively more than 100777Compared to a difference of 15 used in Sec. 2.4. In this simulation, we fit a perfect model (two discs) to a noisy version of itself, so the model fit is already near optimal, resulting in a high likelihood contrast with the case of fitting an imperfect model (single disc). Consequently, we need to adopt a higher BIC threshold compared to fitting models to real data, where both models are imperfect representations of reality. In the real case, unaccounted structures in the data contribute to the $\chi^{2}$, reducing the likelihood contrast between the models.. We visually inspect the fitting results to confirm that the two-discs are identifiable with clarity comparable to the visual inspection process described in Sec. 2.4.
@@ 3.4 The mass ratio of thin and thick discs
In this section, we explore the mass fraction of thick and thin discs in galaxies that host two disc structures. The mass ratio is calculated multiplying the luminosity ratio with the ratio of mass-to-light ratios for thick and thin disks, $M_{\mathrm{thick}}/M_{\mathrm{thin}}=\frac{\Upsilon_{\mathrm{thick}}}{\Upsilon_{\mathrm{thin}}}\frac{L_{\mathrm{thick}}}{L_{\mathrm{thin}}}$. We adopt the fiducial ratio of thin and thick mass-to-light ratio, $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}=1.2$, derived by Comerón et al. (2011) assuming the star formation history (SFH) of the Milky Way’s thin and thick disks at the F356W band (Pilyugin & Edmunds, 1996; Nykytyuk & Mishenina, 2006, but see additional assumptions made in Comerón et al. 2011). Figure 10 shows the mass ratio of thick to thin discs for our sample, as a function of the total stellar mass, compared with $z=0$ galaxies from Comerón et al. (2014) using Spitzer $µm3.5\micron$ band (equivalent to F356W used in this study).
@@
The thick and thin disc mass ratios $M_{\mathrm{thick}}/M_{\mathrm{thin}}$ for our two disc galaxies at $z\sim 0.1-2$ show a decreasing trend as a function of stellar mass. This result is well aligned with massive galaxies at $z=0$ (Comerón et al., 2014). Yoachim & Dalcanton (2006) also derived the mass ratio for their $z=0$ sample using the $R$ band, where the mass-to-light ratios are strongly influenced by galaxies’ SFH. Therefore, we do not make a comparison, but a similar decreasing trend is observed.
To further explore what drives the decreasing trend of thick-to-thin disc luminosity with stellar mass, Fig. 11 shows the individual disc mass of thin and thick discs plotted against total stellar mass. The best-fit slopes $M_{*}-M_{\mathrm{disc}}$ for thin and thick discs are shown. Fits are derived from the two best-fit relations $M_{*}-z_{0}$ and $M_{\mathrm{disc}}-z_{0}$ (Table 1), demonstrating that a single power-law can describe the thin and thick disc sequences separately. The thin and thick disc sequences are consistent with the local results (Comerón et al., 2014), showing two clear distinct sequences with a shallow slope for $M_{\mathrm{thick}}-M_{*}$ and a steep slope for $M_{\mathrm{thin}}-M_{*}$ that cross at a $\log(M_{*}[M_{\odot}])\sim 10$. The different slopes of the thin and thick disc sequences are responsible for the decreasing disc mass ratio as a function of galaxy masses seen in Fig. 10.
@@ 4 Discussion
We confirm the presence of single discs up to $z\sim~{}3$ and two discs up to $z\sim 2$, which already exhibit radial and vertical size-mass relations (Sec.3.1). The edge-on confirmation of stellar discs complements the identification of spiral structures in face-on stellar discs (Kuhn et al., 2024) at redshift up to $z\sim 3$. Well developed stellar discs are internally unstable or dynamically responsive to external perturbations, forming spiral patterns (Byrd & Howard, 1992; Law et al., 2012; Pettitt et al., 2016; Bland-Hawthorn et al., 2023; Tsukui et al., 2024), where the spiral inducing mechanism is presumably important in an increasingly chaotic high redshift universe (Rodriguez-Gomez et al., 2015).
@@ 4.1 Toomre Q self-regulated disc formation
The kinematics of gaseous discs have been systematically characterized up to redshift $z\sim 2.7$ (e.g., Wisnioski et al., 2015; Übler et al., 2019), with more recent studies extending this exploration to redshifts beyond $z\sim 4$ (e.g., Neeleman et al., 2020; Rizzo et al., 2020; Lelli et al., 2021; Tsukui & Iguchi, 2021). The commonly measured kinematic parameter, $v$/$\sigma$, serves as a proxy for the dynamical support of gas discs and provides insights into the geometric proportion of the resulting stellar disc formed from star formation in the gas disc. Fig. 12 shows the intrinsic vertical to radial ‘axial ratio’ of the discs, $h_{R}/z_{0}$, for all disc categories as a function of total stellar mass, as shown in Fig. 5 (right). We compare these measured geometrical proportions with $v/\sigma$ values from the literature for galaxies at similar redshifts $z\sim 0.1-2.7$, mainly measured using optical emission lines (Übler et al., 2019), and for galaxies at earlier redshifts ($>2$) primarily using FIR emission lines (e.g., [Cii], [Ci]; sources in the footnote [^8] Neeleman et al. 2020; Rizzo et al. 2020, 2021; Lelli et al. 2021; Tsukui & Iguchi 2021; Fraternali et al. 2021; Rizzo et al. 2023; Roman-Oliveira et al. 2023; Parlanti et al. 2023; Amvrosiadis et al. 2023; Rowland et al. 2024; Fujimoto et al. 2024). Although measurements by optical emission lines outnumber those using FIR emission measurements, both distributions appear similar when plotted separately in the diagram. Assuming stellar components inherit the same $v/\sigma$ of the gas disc they form out of, a simple application of the tensor virial theorem (Binney, 2005; Binney & Tremaine, 2008) predicts a relation between $h_{R}/z_{0}$ and $v/\sigma$ of self-gravitating axisymmetric stellar structure. Under the assumption of an isotropic velocity dispersion, the approximate relation is given by Kormendy (1982) as,
@@
Gas discs exhibit a weak correlation between $v/\sigma$ (or the expected $h_{R}/z_{0}$) and stellar mass, with more massive galaxies typically located in the thin disc regime and less massive galaxies in the thick disc regime or below. This trend can be explained by self-regulated star formation of gas discs, supporting the downsizing thin disc formation discussed earlier. The Toomre stability criterion for gas-dominated discs (Toomre, 1964) is:
@@
The expected $h_{R}/z_{0}$ ratio for gas discs shows a broader range than observed in stellar discs, indicating that gas discs are subject to significant temporal variation of the Toomre $Q_{\mathrm{gas}}$ parameter. These fluctuations in $Q_{\mathrm{gas}}$ are driven by episodic events such as gas accretion and mergers, which increase the gas fraction ($f_{\mathrm{gas}}$), and subsequent starburst that enhance turbulent energy ($\sigma$), and its dissipation (see e.g., Tacchella et al., 2016). As a result, the $Q_{\mathrm{gas}}$ parameter for gas discs can vary widely, oscillating around the marginal stable value ($Q_{\mathrm{crit}}$). Also, $Q_{\mathrm{crit}}$ itself is variable; it can decrease to about $0.7$ when the finite scale height stabilizes a disc Kim et al. 2002), but can increase to $2-3$ under a condition that gas turbulence efficiently dissipates (Elmegreen 2011). In Fig. 12, we show the variations in $v/\sigma$ expected from the averaged population by assuming either a 2x higher or lower value for $f_{\mathrm{gas}}$ or $Q_{\mathrm{crit}}$, or the product $f_{\mathrm{gas}}Q_{\mathrm{crit}}$, which encompasses the range of gas disc measurements.
In contrast, stellar distributions are shaped by the cumulative effect of star formation, occurring under the variable conditions of the gas disc, presumably leading to a convergence toward narrower axial ratios over time. As the stellar component develops in gas discs, the total disc Toomre parameter, $Q_{\mathrm{tot}}^{-1}=Q_{\mathrm{*}}^{-1}+Q_{\mathrm{gas}}^{-1}$, is subject to the disc stability criterion. This is valid if all components have similar velocity dispersions (Romeo & Wiegert, 2011). Considering a simple scenario, where the gas distribution and stellar distribution are coupled with the same velocity structure and distribution and $f_{\mathrm{gas}}+f_{\mathrm{star}}=1$, the marginally unstable disc would have $h_{R}/z_{0}\sim(v/\sigma)^{2}=(a/Q_{\mathrm{crit}})^{2}$, which does not depend on galaxy stellar mass, consistent with no correlation seen in $h_{R}/z$ for stellar discs in Fig. 12.
Note also that during or after their formation from gaseous discs, stellar discs undergo distinct processes from gaseous discs such as heating by disc substructure or mergers. Unlike gaseous discs, once heated, the stellar components do not cool.
@@ 4.2 Vertical equilibrium of the disc
In the previous section (Sec.4.1), we demonstrated that a gas disc with widely varying gas fractions $f_{\mathrm{gas}}$, can form stellar thin and thick discs with similar geometric proportions -- the relative height to the radial length $h_{R}/z_{0}$. However, this does not necessarily mean that turbulent gaseous discs are sufficiently thick or thin to directly produce thick or thin stellar discs. In this section, we address this question by examining the vertical equilibrium of these system. Figure 13 shows the expected velocity dispersion of the stellar disc assuming vertical equilibrium. The scale height $z_{0}$ of a self-gravitating isothermal sheet is given by $z_{0}=\sigma_{*}^{2}/(2\pi G\Sigma{*})$, where $\sigma_{*}$ is the stellar velocity dispersion, and $\Sigma{*}$ is the surface density of the stellar disc. Using measurements of the scale height $z_{0}$ and the urface density $\Sigma{*}=M_{*}/(2\pi h_{R}^{2})$, we compute the expected velocity dispersion for a single disc in vertical equilibrium as:
@@
Figure 13: Expected stellar velocity dispersion from vertical equilibrium for single, thin and thick discs, with black circles, blue triangles, and red squares, respectively Green circles indicate stellar velocity dispersion measurements from the DiskMass survey for face-on galaxies at $z=0$ (Martinsson et al., 2013b, a). These stellar velocity dispersions are compared with those of gas discs.
For simplicity we exclude contributions from dark matter and small bulges (median contribution of 2% of luminosity). Their inclusion would marginally increase the estimated $\sigma$. The expected velocity dispersion of thin discs aligns with the measurements of vertical stellar velocity dispersion by the DiskMass Survey for local galaxies (Martinsson et al., 2013b, a). The light weighted dispersion of the literature sample traces the thin disc stellar dispersion. This alignment suggests that discs achieve vertical equilibrium, and their measured thickness reflects the virialized state. The vertical equilibrium can be reached rapidly: for instance, the crossing time of a stellar particle within a typical vertical scale height of 0.5 kpc with a velocity dispersion of 50 km s-1 is approximately 10 Myr.
@@
4.3 Adiabatic expansion of the stellar discs
The Milky Way’s thin disc is more radially extended than its thick disc. This contrasts with our measurements, showing the thick disc is often larger than the thin disc, which is also shown in local galaxy samples (Yoachim & Dalcanton, 2006; Comerón et al., 2012).

### v2

@@ 3.3 Emerging thin discs at later epochs
Figure 9 shows the distribution of two-disc and single-disc galaxies in stellar mass and lookback time, along with fractions of two-disc galaxies as a function of each variable. Two-disc galaxies are prevalently found at high stellar masses at earlier times, extending back to $\sim$10 Gyr ago ($z\sim 2$). At later times (i.e., shorter lookback times), they are increasingly found at lower stellar masses, resulting in a increasing fraction of two-disc galaxies with cosmic time and stellar mass. The distribution includes observational biases: identifying a two-disc structure becomes more difficult for more distant galaxies or lower mass galaxies. To dissect observational effects from the onset of thin disc formation, Fig. 9 shows the range of lookback times at which two discs can be identified when a galaxy (of similar properties) with two discs is artificially redshifted (gray line for each data point).
To find the maximum lookback time, we calculate redshifted surface brightness, disc radius ($h_{R}$), and disc height ($z_{0}$) by redshift increments ($dz=0.02$) using galaxies’ best-fit model parameters for galaxies categorised as having two discs (Sec. 3). We consider the surface brightness evolution including not only the surface brightness dimming ($\propto(1+z)^{-4}$) but also k-correction and intrinsic evolution (see Fig. 25 in Appendix E). We then generate simulated images with noise added and the identical PSF, and refit the images using both a two-disc model and one-disc model. The noise incorporates variance from sky noise, read noise, and Poisson variance due to photon counting from the object itself. Bulge components are subject to redshifting but fixed during the refit. We consider the two-discs to be detectable if the BIC difference between the two-disc fit and one-disc fit is conservatively more than 100888Compared to a difference of 15 used in Sec. 2.4. In this simulation, we fit a perfect model (two discs) to a noisy version of itself, so the model fit is already near optimal, resulting in a high likelihood contrast with the case of fitting an imperfect model (single disc). Consequently, we need to adopt a higher BIC threshold compared to fitting models to real data, where both models are imperfect representations of reality. In the real case, unaccounted structures in the data contribute to the $\chi^{2}$, reducing the likelihood contrast between the models.. We visually inspect the fitting results to confirm that the two-discs are identifiable with clarity comparable to the visual inspection process described in Sec. 2.4.
@@ 3.4 The mass ratio of thin and thick discs
In this section, we explore the mass fraction of thick and thin discs in galaxies that host two disc structures. The mass ratio is calculated multiplying the luminosity ratio with the ratio of mass-to-light ratios for thick and thin discs, $M_{\mathrm{thick}}/M_{\mathrm{thin}}=\frac{\Upsilon_{\mathrm{thick}}}{\Upsilon_{\mathrm{thin}}}\frac{L_{\mathrm{thick}}}{L_{\mathrm{thin}}}$. We adopt the fiducial ratio of thin and thick mass-to-light ratio, $\Upsilon_{\mathrm{thick}}/\Upsilon_{\mathrm{thin}}=1.2$, derived by Comerón et al. (2011) assuming the star formation history (SFH) of the Milky Way’s thin and thick discs at the F356W band (Pilyugin & Edmunds, 1996; Nykytyuk & Mishenina, 2006, but see additional assumptions made in Comerón et al. 2011). Figure 10 shows the mass ratio of thick to thin discs for our sample, as a function of the total stellar mass, compared with $z=0$ galaxies from Comerón et al. (2014) using Spitzer $µm3.5\micron$ band (equivalent to F356W used in this study).
@@
The thick and thin disc mass ratios $M_{\mathrm{thick}}/M_{\mathrm{thin}}$ for our two disc galaxies at $z\sim 0.1-2$ show a decreasing trend as a function of stellar mass. This result is well aligned with massive galaxies at $z=0$ (Yoachim & Dalcanton, 2006; Comerón et al., 2011, 2012, 2014). Yoachim & Dalcanton (2006) also derived the mass ratio for their $z=0$ sample using the $R$ band, where the mass-to-light ratios are strongly influenced by galaxies’ SFH. Therefore, we do not make a comparison, but a similar decreasing trend is observed.
To further explore what drives the decreasing trend of thick-to-thin disc luminosity with stellar mass, Fig. 11 shows the individual disc mass of thin and thick discs plotted against total stellar mass. The best-fit slopes $M_{*}-M_{\mathrm{disc}}$ for thin and thick discs are shown. Fits are derived from the two best-fit relations $M_{*}-z_{0}$ and $M_{\mathrm{disc}}-z_{0}$ (Table 1), demonstrating that a single power-law can describe the thin and thick disc sequences separately. The thin and thick disc sequences are consistent with the local results (Comerón et al., 2018), showing two clear distinct sequences with a shallow slope for $M_{\mathrm{thick}}-M_{*}$ and a steep slope for $M_{\mathrm{thin}}-M_{*}$ that cross at a $\log(M_{*}[M_{\odot}])\sim 10$. The different slopes of the thin and thick disc sequences are responsible for the decreasing disc mass ratio as a function of galaxy masses seen in Fig. 10.
@@ 4 Discussion
We confirm the presence of single discs up to $z\sim~{}3$ and two discs up to $z\sim 2$, which already exhibit radial and vertical size-mass relations (Sec.3.1). The edge-on confirmation of stellar discs complements the identification of spiral structures in face-on stellar discs (Kuhn et al., 2024) at redshift up to $z\sim 3$. Well developed stellar discs are internally unstable or dynamically responsive to external perturbations, forming spiral patterns (Byrd & Howard, 1992; Law et al., 2012; Pettitt et al., 2016; Bland-Hawthorn et al., 2023; Tsukui et al., 2024). Spiral-inducing mechanisms driven by external perturbations are presumably more significant at higher redshifts, where merger rates are elevated (Rodriguez-Gomez et al., 2015).
@@ 4.1 Toomre Q self-regulated disc formation
The kinematics of gaseous discs have been systematically characterized up to redshift $z\sim 2.7$ (e.g., Wisnioski et al., 2015; Übler et al., 2019), with more recent studies extending this exploration to redshifts beyond $z\sim 4$ (e.g., Neeleman et al., 2020; Rizzo et al., 2020; Lelli et al., 2021; Tsukui & Iguchi, 2021). The commonly measured kinematic parameter, $v$/$\sigma$, serves as a proxy for the dynamical support of gas discs and provides insights into the geometric proportion of the resulting stellar disc formed from star formation in the gas disc. Fig. 12 shows the intrinsic vertical to radial ‘axial ratio’ of the discs, $h_{R}/z_{0}$, for all disc categories as a function of total stellar mass, as shown in Fig. 5 (right). We compare these measured geometrical proportions with $v/\sigma$ values from the literature for galaxies at similar redshifts $z\sim 0.1-2.7$, mainly measured using optical emission lines (Übler et al., 2019), and for galaxies at earlier redshifts ($>2$) primarily using far-infrared (FIR) emission lines (e.g., [Cii], [Ci]; sources in the footnote [^9] Neeleman et al. 2020; Rizzo et al. 2020, 2021; Lelli et al. 2021; Tsukui & Iguchi 2021; Fraternali et al. 2021; Rizzo et al. 2023; Roman-Oliveira et al. 2023; Parlanti et al. 2023; Amvrosiadis et al. 2023; Rowland et al. 2024; Fujimoto et al. 2024). Although measurements by optical emission lines outnumber those using FIR emission measurements, both distributions appear similar when plotted separately in the diagram.
Assuming stellar components inherit the same $v/\sigma$ of the gas disc they form out of, a simple application of the tensor virial theorem (Binney, 2005; Binney & Tremaine, 2008) predicts a relation between $h_{R}/z_{0}$ and $v/\sigma$ of self-gravitating axisymmetric stellar structure. Under the assumption of an isotropic velocity dispersion, the approximate relation is given by Kormendy (1982) as,
@@
Gas discs exhibit a weak correlation between $v/\sigma$ (or the expected $h_{R}/z_{0}$) and stellar mass, with more massive galaxies typically located in the thin disc regime and less massive galaxies in the thick disc regime or below. This trend is consistent with results from simulations (Pillepich et al., 2019; Kohandel et al., 2024) and can be explained by self-regulated star formation of gas discs, supporting the downsizing thin disc formation discussed earlier. The Toomre stability criterion for gas-dominated discs (Toomre, 1964) is:
@@
The expected $h_{R}/z_{0}$ ratio for gas discs shows a broader range than observed in stellar discs, indicating that gas discs are subject to significant temporal variation of the Toomre $Q_{\mathrm{gas}}$ parameter. These fluctuations in $Q_{\mathrm{gas}}$ are driven by episodic events such as gas accretion and mergers, which increase the gas fraction ($f_{\mathrm{gas}}$), and subsequent starburst that enhance turbulent energy ($\sigma$), and its dissipation (see e.g., Tacchella et al., 2016). As a result, the $Q_{\mathrm{gas}}$ parameter for gas discs can vary widely, oscillating around the marginal stable value ($Q_{\mathrm{crit}}$). Also, $Q_{\mathrm{crit}}$ itself is variable; it can decrease to about $0.7$ when the finite scale height stabilizes a disc (Kim et al., 2002; Bacchini et al., 2024, as opposed to an infinitesimally thin disc), but can increase to $2-3$ under a condition that gas turbulence efficiently dissipates (Elmegreen 2011). In Fig. 12, we show the variations in $v/\sigma$ expected from the averaged population by assuming either a 2x higher or lower value for $f_{\mathrm{gas}}$ or $Q_{\mathrm{crit}}$, or the product $f_{\mathrm{gas}}Q_{\mathrm{crit}}$, which encompasses the range of gas disc measurements.
In contrast, stellar distributions are shaped by the cumulative effect of star formation, occurring under the variable conditions of the gas disc, presumably leading to a convergence toward a narrower range of axial ratios over time. As the stellar component develops in gas discs, the total disc Toomre parameter, $Q_{\mathrm{tot}}^{-1}=Q_{\mathrm{*}}^{-1}+Q_{\mathrm{gas}}^{-1}$, is subject to the disc stability criterion. This is valid if all components have similar velocity dispersions (Romeo & Wiegert, 2011). Considering a simple scenario, where the gas distribution and stellar distribution are coupled with the same velocity structure and distribution and $f_{\mathrm{gas}}+f_{\mathrm{star}}=1$, the marginally unstable disc would have $h_{R}/z_{0}\sim(v/\sigma)^{2}=(a/Q_{\mathrm{crit}})^{2}$, which does not depend on galaxy stellar mass, consistent with no correlation seen in $h_{R}/z$ for stellar discs in Fig. 12. Note also that during or after their formation from gaseous discs, stellar discs undergo distinct processes from gaseous discs such as heating by disc substructure or mergers. Unlike gaseous discs, once heated, the stellar components do not cool.
Finally, note that stellar masses of galaxies with available optical line kinematics, derived using the same methods (3D-HST; Übler et al. 2019), are larger than those in our sample. This reflects the current sensitivity limitations of spectroscopic observations and the challenge of obtaining kinematics for low-mass galaxies.
@@ 4.2 Vertical equilibrium of the disc
In the previous section (Sec.4.1), we demonstrated that a gas disc with widely varying gas fractions $f_{\mathrm{gas}}$, can form stellar thin and thick discs with similar geometric proportions -- the relative height to the radial length $h_{R}/z_{0}$. However, this does not necessarily mean that turbulent gaseous discs are sufficiently thick or thin to directly produce thick or thin stellar discs. In this section, we address this question by examining the vertical equilibrium of these system. Figure 13 shows the expected velocity dispersion of the stellar disc assuming vertical equilibrium. The scale height $z_{0}$ of a self-gravitating isothermal sheet is given by $z_{0}=\sigma_{*}^{2}/(2\pi G\Sigma{*})$ (Binney & Tremaine, 2008), where $\sigma_{*}$ is the stellar velocity dispersion, and $\Sigma{*}$ is the surface density of the stellar disc. Using measurements of the scale height $z_{0}$ and the surface density $\Sigma{*}=M_{*}/(2\pi h_{R}^{2})$, we compute the expected velocity dispersion for a single disc in vertical equilibrium as:
@@
Figure 13: Expected stellar velocity dispersion from vertical equilibrium for single, thin and thick discs, with black circles, blue triangles, and red squares, respectively. The arrows indicate upper limit measurements. Green circles indicate stellar velocity dispersion measurements from the DiskMass survey for face-on galaxies at $z=0$ (Martinsson et al., 2013b, a). These stellar velocity dispersions are compared with those of gas discs (black dots).
For simplicity, we exclude contributions from dark matter, as baryons dominate in the disc mid-plane (Narayan & Jog, 2002), and neglect small bulges, which contribute a median of 2% to the total disc luminosity in our sample. Their inclusion would marginally increase the estimated $\sigma$. The expected velocity dispersion of thin discs aligns with the measurements of vertical stellar velocity dispersion by the DiskMass Survey for local galaxies (Martinsson et al., 2013b, a). The light weighted dispersion of the literature sample traces the thin disc stellar dispersion. This alignment suggests that discs achieve vertical equilibrium, and their measured thickness reflects the virialized state. The vertical equilibrium can be reached rapidly: for instance, the crossing time of a stellar particle within a typical vertical scale height of 0.5 kpc with a velocity dispersion of 50 km s-1 is approximately 10 Myr.
@@
4.3 On why thin discs are smaller than thick discs
The Milky Way’s thin disc is more radially extended than its thick disc. This contrasts with our measurements, showing the thick disc is often larger than the thin disc, which is also shown in local galaxy samples structurally (Yoachim & Dalcanton, 2006; Comerón et al., 2012) and chemically (Sattler et al., 2023, 2024).